# WiMANS: A Benchmark Dataset for WiFi-based Multi-user Activity Sensing

### 24 Jan 2024 v1

@@ Abstract
WiFi-based human sensing has exhibited remarkable potential to analyze user behaviors in a non-intrusive and device-free manner, benefiting applications as diverse as smart homes and healthcare. However, most previous works focus on single-user sensing, which has limited practicability in scenarios involving multiple users. Although recent studies have begun to investigate WiFi-based multi-user activity sensing, there remains a lack of benchmark datasets to facilitate reproducible and comparable research. To bridge this gap, we present WiMANS, to our knowledge, the first dataset for multi-user activity sensing based on WiFi. WiMANS contains over 9.4 hours of WiFi Channel State Information (CSI), monitoring simultaneous activities performed by multiple users in various environments. Compared to existing datasets, WiMANS not only collects the CSI of dual WiFi bands but also includes synchronized videos. We exploit WiMANS to benchmark the performance of state-of-the-art WiFi-based human sensing models and video-based models, posing new challenges and opportunities for WiFi-based multi-user identification, localization, and activity recognition. We believe that WiMANS can push the boundaries of current WiFi-based human sensing and catalyze the research on multi-user activity analysis.
@@ 1 Introduction
per Sample Samples Activities Channels Rate (Hz) (GHz) Data Identity Location Activity
@@
Table 1: Comparison of public WiFi-based human sensing datasets. WiMANS is the first dataset that includes multiple users per sample.
@@
Recent years have witnessed the rapid progress of WiFi-based human sensing [41, 22, 5], which collects Channel State Information (CSI) from off-the-shelf WiFi devices to recognize human identities [57, 47, 62], locations [40, 38, 8], activities [28, 17, 6], etc. It plays an increasingly important role in differing applications such as security monitoring [42, 30, 68], smart homes [9, 24], and healthcare [59, 15, 36]. Compared with cameras and on-body sensors, the use of WiFi CSI negates the necessity of filming users or attaching sensors to them [35]. Such non-intrusive approaches can make human sensing widely available [67], satisfying the need to monitor users who do not want to be filmed or wear sensors. WiFi-based human sensing is also robust to low-light or non-line-of-sight conditions, while using cameras for video-based analysis may be sensitive to low-light conditions and obstacles [27]. More importantly, CSI can be gathered from ubiquitous existing WiFi devices, enabling device-free human sensing [25] without the prerequisite of dedicated devices.
Figure 1: WiMANS collects WiFi CSI and videos to monitor simultaneous activities of multiple users. Human activities interfere with WiFi signals via reflection and/or refraction, manifesting distinct WiFi CSI patterns with implicit features for human sensing.
The principle of WiFi-based human sensing is that human activities essentially interfere with WiFi signals and lead to signal variations [51]. Such variations are recorded in CSI, which thereby contains implicit features for human sensing, as shown in Figure 1. Since these human features typically intertwine with excessive noise [50], it is impossible to interpret WiFi CSI patterns for human sensing straightforwardly. Therefore, various models have been proposed to learn features from the CSI for different human sensing tasks. (1) For human identification, extensive literature has discussed the use of Multilayer Perceptrons (MLPs) [58], Long Short-Term Memory (LSTM) [7], Convolutional Neural Networks (CNNs) [47, 57, 56], CNN-LSTM hybrids [27, 62, 33], etc. (2) For human localization, prevalent models have been adopted, including Naive Bayes [52], Auto-encoders [14], LSTM [8], and CNNs [49]. (3) WiFi-based human activity recognition has drawn the greatest attention from researchers, applying models such as MLPs [61], LSTM [60], CNNs [34, 65, 63], CNN-LSTM hybrids [69, 19, 39], Generative Adversarial Networks (GANs) [46], attention-based bidirectional LSTM (ABLSTM) [6], and Transformers [66, 28]. However, several critical drawbacks seriously hamper the practicability of WiFi-based human sensing. (1) Single-user Limitation: Existing public datasets only include a single user in each CSI sample, as shown in Table 1. Consequently, most WiFi-based models only recognize the identity/location/activity of a single user in each recognition, but many practical scenarios indeed consist of multiple users simultaneously. Although several recent studies [21, 40, 45, 11] have attempted to sense multiple users, there is a lack of public datasets that enable WiFi-based multi-user sensing. (2) Insufficient Modalities and Annotations: Most datasets collect CSI of a single WiFi band (i.e., 2.4 or 5 GHz) and do not incorporate synchronized videos. Such insufficiency disables the further study of unexplored tasks (e.g., pose estimation). Moreover, many datasets annotate samples for specific tasks (e.g., activity recognition), restricting the use of WiFi CSI in diverse sensing tasks. (3) Lack of Comprehensive Benchmarks: Previous works mainly focus on novel models but few of them have provided benchmarks for these models. Specifically, no benchmark is available for WiFi-based multi-user sensing. To help resolve the above drawbacks, we present WiMANS, the first dataset that enables multi-user sensing based on WiFi CSI. WiMANS collects 11286 CSI samples of dual WiFi bands (2.4 / 5 GHz), along with synchronized videos for reference, as shown in Figure 2. Each 3-second sample includes 0 to 5 users performing identical/different activities simultaneously, annotated with (anonymized) user identities, locations, and activities. Table 1 compares WiMANS with public WiFi-based human sensing datasets to highlight our novelty. The main contributions and unique aspects of WiMANS are as follows:
  2. WiMANS provides fine-grained annotations of user identities, locations, and activities to support various sensing tasks. The videos in WiMANS can further act as a reference for unexplored tasks (e.g., pose estimation).
@@
Figure 2: Examples of WiFi CSI and videos in WiMANS, monitoring simultaneous activities performed by different numbers of users in various environments. (a) consists of 3 users in a classroom. (b) contains 4 users in a meeting room. (c) includes 5 users in an empty room.
@@ 2.1 WiFi-based Human Sensing
Human sensing with WiFi CSI is an increasingly promising alternative to traditional sensing technologies thanks to its non-intrusiveness, environmental robustness, and device-free merits. In WiFi-based human sensing, much effort has been devoted to three underlying yet distinct tasks: human identification, localization, and activity recognition (HAR). (1) WiFi-based human identification serves to determine user identities by learning biometric patterns from CSI. To simply map CSI to identities, MLPs [58] feed all CSI values into fully connected layers, where excessive parameters lead to high complexity and poor generalization. In contrast, LSTM [7] regards CSI as sequences to learn temporal features. Despite the advanced performance than MLPs, LSTM is inefficient for lengthy sequences due to its step-by-step inputs. To boost both effectiveness and efficiency, CNNs [47, 57, 56] have been utilized by dividing the CSI into diverse receptive fields to learn spatial features. Finally, CNN-LSTM hybrids [27, 62, 33] have been proposed to exert the strengths of CNNs and LSTM, attaining the best performance for WiFi-based human identification. (2) WiFi-based human localization estimates user locations to support human-computer interaction systems [41]. Initially, Naive Bayes [52] has been leveraged for localization, showing acceptable performance. Thereafter, Sparse Auto-encoder (SAE) [14] has been used to localize users based on CSI images, but its fully connected layers still lead to excessive parameters. LSTM [8] and CNNs [49] have been further explored for localization, where CNN-1D [49] shows its superior capability. (3) WiFi-based HAR is gaining popularity for user behavior analysis. Initial methods [60] based on handcrafted feature extraction have demonstrated inadequate ability to extract implicit features from CSI. Thereby, LSTM [60] and CNNs [34, 63, 65] have been extensively adopted and offer valuable insights on learning temporal and spatial features. CNN-LSTM hybrids [69, 19, 39] further become predominant models for HAR. To tackle dynamic environments, GANs [46] have been used to augment CNNs with adversarial learning. Attention-based models [44] also contribute to HAR [66], such as ABLSTM [6] which equips bidirectional LSTM with attention layers to enhance HAR performance. Recently, a two-stream convolution augmented transformer (THAT) [28] has further integrated attention layers with multi-scale CNNs, achieving state-of-the-art performance. However, all the above methods carry out single-user sensing only, while multi-user sensing based on WiFi CSI is more practical yet challenging. Recent works [21, 40, 45, 11] have begun to sense multiple users simultaneously with their own CSI samples, but they have not published their datasets. Such an obvious need motivates WiMANS as the first dataset that involves multiple users in each CSI sample.
@@
Figure 3: Layouts of environments in WiMANS, where a transmitter and a receiver record WiFi CSI, and a monitor camera captures videos.
2.2 Datasets for WiFi-based Human Sensing
High-quality annotated CSI datasets are essential for advancing WiFi-based human sensing research. Table 1 summarizes existing CSI datasets in comparison to WiMANS. At the beginning, Yousefi et al. [60] collected a dataset in an office for WiFi-based HAR, after which various CSI datasets were released for HAR [3, 18, 55, 13]. Baha et al. [1] further constructed a dataset under line-of-sight and non-line-of-sight conditions. Some datasets were presented for specific tasks, such as SignFi [32] for sign language recognition, FallDeFi [36] for fall detection, and ARIL [48] for activity recognition and localization. Widar [37] and Widar 2.0 [38] were gathered for indoor human tracking. Some datasets contain changing scenarios, including CPAR [54] for different users and RF-NET [10] for varying environments. Similarly, CSIDA [23] and Widar 3.0 [64] incorporate CSI of various domains (i.e., locations, orientations, environments, and users). OPERAnet [2], a multi-modal dataset, includes WiFi CSI as well as data from Passive WiFi Radar (PWR), Ultra-Wideband (UWB), and Kinect skeleton sensors. NTU-Fi [58] further uses its WiFi CSI dataset to benchmark human sensing models. Despite these datasets, none of them enables simultaneous multi-user sensing, which is crucial for practical WiFi-based human sensing advancement, since real-life scenarios typically involve multiple users simultaneously. To bridge this gap, we propose WiMANS, the first dataset for WiFi-based multi-user activity sensing, hoping to facilitate the next generation of human sensing based on WiFi.
3 The WiMANS Dataset
WiMANS aims to gather WiFi CSI samples which monitor simultaneous activities of multiple users. Each CSI sample contains 0 to 5 users performing identical/different activities at the same time, as shown in Figure 2. Annotations are provided for all samples, including user identities, locations, and activities. Ultimately, WiMANS collects 11286 samples (over 9.4 hours) of WiFi CSI and videos.
3.1 Dataset Construction
WiFi CSI monitors the variations of wireless signals by which a transmitter propagates packets to a receiver. To collect WiFi CSI, we deploy two off-the-shelf computers (HP EliteDesk 800 G2 TWR) to serve as the transmitter and receiver. Each device is equipped with an Intel 5300 Network Interface Card and has the Linux 802.11n CSI tool [20] installed, following the previous works [60, 64, 2]. We set these devices to work in the monitor mode, enabling us to control the start and the end of each CSI sample. To exert the advantages of different WiFi bands, we collect dual-band WiFi CSI by setting devices to work on channel 12 for the 2.4 GHz band, and on channel 64 for the 5 GHz band. Data Collection. We employ the transmitter and receiver to collect a CSI sample in 3 steps: (1) The receiver listens to a WiFi channel and logs the CSI from all packets it receives; (2) The transmitter sends packets to the WiFi channel, and the users simultaneously perform designated activities; (3) The receiver stops logging and listening. Following the previous works [64, 2], we instruct users to perform activities in 3 seconds and thereby control the transmitter to send 3000 packets at a rate of 1000 packets per second. Regardless of packet loss, each 3-second CSI sample should have 3000 time steps. We will further analyze the issue of packet loss in Section 3.2. The transmitter and receiver individually have 3 antennas, and each pair of antennas uses 30 subcarriers for wireless communication. Thus, the dimension of CSI at each time step is 3$\times$3$\times$30, and the dimension of each CSI sample is 3000$\times$3$\times$3$\times$30. Along with each CSI sample, we also capture synchronized videos with a monitor camera for reference and unexplored tasks (e.g., pose estimation). Each 3-second video has 90 frames under the frame rate of 30 Hz, with 3 RGB channels and the frame resolution of 1920$\times$1080. Hence, the dimension of each video sample is 90$\times$3$\times$1920$\times$1080. Collection Setup. WiMANS includes 9 activities which are representative in daily life, as discussed in the previous works [60, 64, 2]. These activities are: (1) Nothing, (2) Walking, (3) Rotation, (4) Jumping, (5) Waving, (6) Lying Down, (7) Picking Up, (8) Sitting Down, (9) Standing Up. To build up a comprehensive dataset, we collect data in 3 daily environments: (1) Classroom, (2) Meeting Room, (3) Empty Room. In each environment, we specify 5 locations and mark them as A, B, C, D, and E. Figure 3 describes the layouts of environments and the interior locations. We recruit 6 volunteers to act as users, including 3 females and 3 males, with an average age of 27.33 $\pm$ 0.94, height of 169.00 $\pm$ 6.83 cm, weight of 61.17 $\pm$ 10.04 kg, and Body Mass Index (BMI) of 21.25 $\pm$ 1.78. These users are assigned with identity labels (e.g., User 1, User 2, …) to protect their privacy and to support human identification. Regarding the human subject study, we have obtained an approval from institutional ethics committee (IRB) in advance and will provide an ethics statement in Section 5.
Figure 4: Scripts that instruct users to perform activities at different locations independently yet simultaneously.
@@ 3.1 Dataset Construction
We control the collection of each 3-second sample in three dimensions: (1) users, (2) locations, (3) activities. For the user dimension, we organize volunteers into user groups, where each group includes fixed users. For the location and activity dimensions, we devise simultaneous scripts for users in each group, so that they know where to stay and what to do independently yet simultaneously. User Groups. We collect samples by user groups, where each group corresponds to a specific number of users and environment. For example, Figure 1(a) shows a sample in the 30th user group, where 3 users (User 1, 5, and 6) are in the classroom. Specifically, there are 6 groups for 0 user, 36 groups for 1 user, 18 groups for 2 users, 18 groups for 3 users, 18 groups for 4 users, and 18 groups for 5 users. On the other hand, there are 38 groups for each environment. Each group contains 99 samples, and we totally obtain 11286 samples for 114 groups in WiMANS. In each group, we design scripts for users to perform activities. Simultaneous Scripts. In each group, we allocate scripts to fixed users, instructing them to perform activities at different locations independently yet simultaneously. To this end, we devise a script set $\bm{\alpha}$, $\bm{\beta}$, $\bm{\delta}$, $\bm{\gamma}$, $\bm{\lambda}$}, as shown in Figure 4. Each script has 99 indexes, and each index corresponds to an activity at a location. Once we announce an index, users with different scripts realize where to stay and what to do. In Figure 1(a), after we announce the index 25, User 1 with $\bm{\alpha}$ performs Picking Up, User 5 with $\bm{\lambda}$ performs Walking, and User 6 with $\bm{\beta}$ performs Sitting Down. Note that the scripts are put away before activities to minimize their impacts on data collection. Between the collection of each two samples, we first stop the collection of previous sample and then announce a new index to users, after which we start the collection of next sample. Therefore, we can ensure the purity of each sample, not containing nearby activities. More details of activities, users, user groups, and simultaneous scripts are provided in Appendix A. Data Annotation. Since we apply the aforementioned scripts to instruct user activities, we have actually created data annotations before collection. We label each sample as “act_$<$group$>$_$<$sample$>$”, where “$<$group$>$” is the user group index, and “$<$sample$>$” is the sample index. Figure 1(a) shows the sample labeled as “act_30_25”, manifesting that it is the 25th sample of the 30th user group. We associate each label with the environment, WiFi band, number of users, user identities, locations, and activities, to enable different tasks. Note that the videos share the same labels with CSI and thereby can act as references to support unexplored tasks (e.g., WiFi-based multi-user pose estimation).

### 12 Mar 2024 v2

# https://github.com/huangshk/WiMANS

@@ Abstract
WiFi-based human sensing has exhibited remarkable potential to analyze user behaviors in a non-intrusive and device-free manner, benefiting applications as diverse as smart homes and healthcare. However, most previous works focus on single-user sensing, which has limited practicability in scenarios involving multiple users. Although recent studies have begun to investigate WiFi-based multi-user sensing, there remains a lack of benchmark datasets to facilitate reproducible and comparable research. To bridge this gap, we present WiMANS, to our knowledge, the first dataset for multi-user sensing based on WiFi. WiMANS contains over 9.4 hours of dual-band WiFi Channel State Information (CSI), as well as synchronized videos, monitoring simultaneous activities of multiple users. We exploit WiMANS to benchmark the performance of state-of-the-art WiFi-based human sensing models and video-based models, posing new challenges and opportunities for future work. We believe WiMANS can push the boundaries of current studies and catalyze the research on WiFi-based multi-user sensing.
Keywords:
Human Sensing Multi-user WiFi Benchmark Dataset
@@ 1 Introduction
Table 1: Comparison of public WiFi-based human sensing datasets. WiMANS is the first dataset that involves multiple users performing different/identical activities simultaneously in each sample. “Idt.”: Identity. “Loc.”: Location. “Act.”: Activity.
@@
per Sample Samples Activities Channels Rate (Hz) (GHz) Data Idt. Loc. Act.
@@
Recent years have witnessed the rapid progress of WiFi-based human sensing [49, 26, 6], which collects Channel State Information (CSI) from off-the-shelf WiFi devices to recognize human identities [69, 56, 76], locations [12, 46, 9], activities [34, 21, 7], etc. It plays an increasingly important role in differing applications such as security monitoring [51, 37, 84], smart homes [10, 28], and healthcare [72, 16, 44]. Compared with cameras and on-body sensors, the use of WiFi CSI negates the necessity of filming users or attaching sensors to them [43]. Such non-intrusive approaches can make human sensing widely available [82], satisfying the need to monitor users who do not want to be filmed or wear sensors. WiFi-based human sensing is also robust to low-light or non-line-of-sight conditions, while using cameras for video-based analysis may be sensitive to low-light conditions and obstacles [33]. More importantly, CSI can be gathered from ubiquitous existing WiFi devices, enabling device-free human sensing [29] without the prerequisite of dedicated devices and particular deployments. The principle of WiFi-based human sensing is that human activities essentially interfere with WiFi signals and lead to signal variations [62]. Such variations are recorded in CSI, which thereby contains implicit features for human sensing. Since these human features typically intertwine with excessive noise [61], it is impossible to interpret WiFi CSI patterns for human sensing straightforwardly. Therefore, various models have been proposed to learn features from the CSI for different human sensing tasks. (1) For human identification, extensive literature has discussed the use of Multilayer Perceptrons (MLPs) [67], Long Short-Term Memory (LSTM) [8], Convolutional Neural Networks (CNNs) [56, 69, 68], CNN-LSTM hybrids [33, 76, 41], etc. (2) For human localization, prevalent models have been adopted, including Naive Bayes [63], Auto-encoders [15], LSTM [9], and CNNs [58]. (3) WiFi-based human activity recognition (HAR) has drawn the greatest attention from researchers, applying models such as MLPs [74], LSTM [73], CNNs [42, 78, 75], CNN-LSTM hybrids [85, 23, 48], Generative Adversarial Networks (GANs) [55], attention-based bidirectional LSTM (ABLSTM) [7], and Transformers [81, 34].
@@
Figure 1: Examples of WiFi CSI and videos in WiMANS, monitoring simultaneous activities performed by multiple users in various environments. (a) consists of 3 users in a classroom. (b) contains 4 users in a meeting room. (c) has 5 users in an empty room.
However, several critical drawbacks seriously hamper the practicability of WiFi-based human sensing. (1) Single-user Limitation: Existing public datasets only include a single user in each CSI sample, as shown in Table 1. Consequently, most WiFi-based models only recognize the identity/location/activity of a single user in each recognition, but many practical scenarios indeed consist of multiple users simultaneously. Although recent studies [25, 12, 32, 59] have attempted to sense multiple users simultaneously, there remains a lack of public datasets that enable WiFi-based multi-user sensing. (2) Insufficient Modalities and Annotations: Most datasets collect CSI of a single WiFi band (i.e., 2.4 or 5 GHz) and do not incorporate synchronized videos. Such insufficiency disables the further study of unexplored tasks (e.g., pose estimation). Moreover, many datasets annotate samples for specific tasks (e.g., activity recognition), restricting the use of WiFi CSI in diverse sensing tasks. (3) Lack of Comprehensive Benchmarks: Previous works mainly focus on novel models but few of them have provided benchmarks for these models. Specifically, no benchmark is available for multi-user activity sensing based on WiFi CSI. To help resolve the above drawbacks, we present WiMANS, the first dataset that enables multi-user sensing based on WiFi CSI. WiMANS collects 11286 CSI samples of dual WiFi bands (2.4 / 5 GHz), along with synchronized videos for reference, as shown in Fig. 1. Each 3-second sample includes 0 to 5 users performing identical/different activities simultaneously, annotated with (anonymized) user identities, locations, and activities. Table 1 compares WiMANS with related datasets to highlight our novelty. Extensive experiments have been conducted to benchmark state-of-the-art WiFi-based models and video-based models. The main contributions and unique aspects of WiMANS are as follows:
  2. WiMANS provides fine-grained annotations of user identities, locations, and activities to support various sensing tasks. The videos in WiMANS can further act as a reference for unexplored tasks (e.g., multi-user pose estimation).
@@ 2.1 WiFi-based Human Sensing
Human sensing with WiFi CSI is an increasingly promising alternative to traditional sensing technologies thanks to its non-intrusiveness, environmental robustness, and device-free merits. In WiFi-based human sensing, much effort has been devoted to three underlying yet distinct tasks: (1) human identification, (2) human localization, and (3) human activity recognition (HAR). WiFi-based human identification serves to determine user identities by learning biometric patterns from CSI. To map CSI to identities, MLPs [67] simply feed all CSI values into fully connected layers, where excessive parameters lead to high complexity and poor generalization. In contrast, LSTM [8] regards CSI as sequences to learn temporal features. Despite the advanced performance than MLPs, LSTM is inefficient for lengthy sequences due to its step-by-step inputs. To boost both effectiveness and efficiency, CNNs [56, 69, 68] have been utilized by dividing the CSI into diverse receptive fields to learn spatial features with convolution filters. To combine the advantages of CNN and LSTM, recent works [33, 76, 41] have proposed various CNN-LSTM hybrids, attaining the best performance for WiFi-based human identification. WiFi-based human localization estimates user locations to support human-computer interaction systems [49]. Initially, Naive Bayes [63], a basic statistic model, has been leveraged for localization, showing acceptable performance. Thereafter, Sparse Auto-encoder (SAE) [15] has been used to localize users based on CSI images, but its fully connected layers still lead to excessive parameters. LSTM [9] and CNNs [58] have been further explored for localization, where CNN-1D [58] shows its superior capability for WiFi-based human localization. WiFi-based HAR is gaining popularity for user behavior analysis, which is of finer granularity than identification and localization. Initial methods [73] based on handcrafted feature extraction have demonstrated inadequate ability to extract implicit features from CSI. Thereby, LSTM [73] and CNNs [42, 75, 78] have been extensively adopted and offer valuable insights on learning temporal and spatial features. CNN-LSTM hybrids [85, 23, 48] further leverage the advantages of CNNs and LSTM to become predominant models for HAR. To tackle dynamic environments, GANs [55] have been used to augment CNNs with adversarial learning. Attention-based models [53] also contribute to HAR [81], such as ABLSTM [7] which equips bidirectional LSTM with attention layers to enhance HAR performance. Recently, a two-stream convolution augmented transformer (THAT) [34] has further integrated attention layers with multi-scale CNNs, achieving state-of-the-art performance for WiFi-based HAR. However, all the above methods carry out single-user sensing only, while multi-user sensing based on WiFi CSI is more practical yet challenging. Multiple users in an environment may occlude each other, challenging the resolution of models. The mutual interference between users also requires models to disentangle features of different users for effective sensing. Recent works [25, 12, 32, 59, 50, 54] have attempted to tackle these challenges and sense multiple users simultaneously, but they have not published their datasets and have not provided any benchmark. Such an obvious need motivates WiMANS as the first benchmark dataset that involves multiple users in each CSI sample.
2.2 Datasets for WiFi-based Human Sensing
High-quality annotated CSI datasets are essential for advancing WiFi-based human sensing research. Table 1 summarizes existing CSI datasets in comparison to WiMANS. At the beginning, Yousefi et al. [73] collected a dataset in an office for WiFi-based HAR, after which various CSI datasets were released for HAR in different environments [4, 22, 71, 14]. Baha et al. [2] further constructed a dataset under line-of-sight and non-line-of-sight conditions. Some datasets were presented for specific tasks, such as SignFi [39] for sign language recognition, FallDeFi [44] for fall detection, and ARIL [57] for activity recognition and localization. For human tracking, Widar [45] and Widar 2.0 [46] were gathered and annotated with user walking traces. Some datasets contain changing scenarios, such as CPAR [66] for different users and RF-NET [11] for varying environments. Similarly, CSIDA [27] and Widar 3.0 [77] incorporate CSI of various domains (i.e., locations, orientations, environments, and users). OPERAnet [3], a multi-modal dataset, includes WiFi CSI as well as data from Passive WiFi Radar (PWR), Ultra-Wideband (UWB), and Kinect skeleton sensors. MM-Fi [70] is also a multi-modal dataset, containing WiFi CSI, video frames, depth frames, LiDAR point cloud, and mmWave radar point cloud. For benchmarking, Yang et al. [67] collected the NTU-Fi dataset to compare WiFi-based human sensing models in terms of their performance, sizes, complexity, etc. To analyze the impact of wireless parameters (e.g., bandwidth), Meneghello et al. [40] constructed the SHARPax dataset to explore WiFi-based HAR under varying settings. Despite these datasets, none of them enables simultaneous multi-user sensing, which is crucial for practical WiFi-based human sensing advancement, since real-life scenarios typically involve multiple users simultaneously. To bridge this gap, we propose WiMANS, the first dataset for WiFi-based multi-user activity sensing, hoping to facilitate the next generation of human sensing based on WiFi.
3 WiMANS
WiMANS aims to gather WiFi CSI samples which monitor simultaneous activities of multiple users. Each CSI sample contains 0 to 5 users performing identical/different activities at the same time, as shown in Fig. 1. Fine-grained annotations are provided for all samples, including user identities, locations, and activities. Ultimately, WiMANS collects 11286 samples (over 9.4 hours) of dual-band WiFi CSI and synchronized videos.
3.1 Dataset Construction
Hardware Setup. WiFi CSI monitors the variations of wireless signals by which a transmitter propagates packets to a receiver. To collect WiFi CSI, we deploy two off-the-shelf computers (HP EliteDesk 800 G2 TWR) to serve as the transmitter and receiver. Each device is equipped with an Intel 5300 Network Interface Card and has the Linux 802.11n CSI tool [24] installed, following the previous works [73, 77, 3]. We set these devices to work in the monitor mode, enabling us to control the start and the end of each CSI sample. To exert the advantages of different WiFi bands, we collect dual-band WiFi CSI by setting devices to work on channel 12 for the 2.4 GHz band, and on channel 64 for the 5 GHz band. Meanwhile, we use a monitor camera to capture synchronized videos. Data Collection. We employ the transmitter and receiver to collect a CSI sample in 3 steps: (1) The receiver listens to a WiFi channel and logs the CSI from all packets it receives; (2) The transmitter sends packets to the WiFi channel, and the users simultaneously perform designated activities; (3) The receiver stops logging and listening. Following the previous works [77, 3], we instruct users to perform activities in 3 seconds and thereby control the transmitter to send 3000 packets at a rate of 1000 packets per second. Regardless of packet loss, each 3-second CSI sample should have 3000 time steps under the sample rate of 1000 Hz. We will further analyze the issue of packet loss in Section 3.2. The transmitter and receiver individually have 3 antennas, and each pair of antennas uses 30 subcarriers for wireless communication. Thus, the dimension of CSI at each time step is 3$\times$3$\times$30, and the dimension of each CSI sample is 3000$\times$3$\times$3$\times$30. Along with each CSI sample, we also capture synchronized videos with a monitor camera for reference and unexplored tasks (e.g., multi-user pose estimation). Each 3-second video has 90 frames under the frame rate of 30 Hz, with 3 RGB channels and the frame resolution of 1920$\times$1080. Hence, the dimension of each video sample is 90$\times$3$\times$1920$\times$1080.
Data Synchronization. During data collection, WiFi CSI is collected sample by sample, while long videos are recorded at the same time. Afterward, we synchronize WiFi CSI and videos by segmenting long videos into 3-second samples based on timestamps. Because the frame rate of videos is 30 Hz while the sample rate of CSI is 1000 Hz, we can synchronize WiFi CSI and video samples in 16.67 ms, which is unavoidable and tolerable for 3-second samples. Data Attributes. WiMANS includes 9 activities which are representative in daily life, as discussed in the previous works [73, 77, 3]. These activities are: (1) Nothing, (2) Walking, (3) Rotation, (4) Jumping, (5) Waving, (6) Lying Down, (7) Picking Up, (8) Sitting Down, (9) Standing Up. To build up a comprehensive dataset, we collect data in 3 daily environments: (1) Classroom, (2) Meeting Room, (3) Empty Room. In each environment, we specify 5 locations and mark them as A, B, C, D, and E. Fig. 2 describes the layouts of three environments and the interior locations. We recruit 6 volunteers to act as users, including 3 females and 3 males, with an average age of 27.33 $\pm$ 0.94, height of 169.00 $\pm$ 6.83 cm, weight of 61.17 $\pm$ 10.04 kg, and Body Mass Index (BMI) of 21.25 $\pm$ 1.78. These users are assigned with identity labels (e.g., User 1, User 2, …) to protect their privacy and to support human identification. Regarding the human subject study, we have obtained an approval from institutional ethics committee (IRB) in advance and will provide an ethics statement in Section 5.3.
@@ 3.1 Dataset Construction
Figure 2: Layouts of environments in WiMANS, where a transmitter and a receiver collect WiFi CSI, and a monitor camera captures synchronized videos for reference.
@@
Collection Setup. We control the collection of each 3-second sample in three dimensions: (1) users, (2) locations, (3) activities. For the user dimension, we organize volunteers into user groups, where each group includes fixed users. For the location and activity dimensions, we devise simultaneous scripts for users in each group, so that they know where to stay and what to do independently yet simultaneously. User groups and simultaneous scripts are described as follows. User Groups. We collect samples by user groups, where each group corresponds to a specific number of users and environment. For example, Fig. 0(a) shows a sample in the 30th user group, where 3 users (User 1, 5, and 6) are in the classroom. Specifically, there are 6 groups for 0 user, 36 groups for 1 user, 18 groups for 2 users, 18 groups for 3 users, 18 groups for 4 users, and 18 groups for 5 users. On the other hand, there are 38 groups for each environment. Each group contains 99 samples, and we totally obtain 11286 samples for 114 user groups in WiMANS. In each user group, we design simultaneous scripts for users to perform activities. Simultaneous Scripts. In each group, we allocate scripts to fixed users, instructing them to perform activities at different locations independently yet simultaneously. To this end, we devise a script set $\bm{\alpha}$, $\bm{\beta}$, $\bm{\delta}$, $\bm{\gamma}$, $\bm{\lambda}$}, as shown in Fig. 3. Each script has 99 indexes, and each index corresponds to an activity at a location. Once we announce an index, users with different scripts realize where to stay and what to do. For example, in Fig. 0(a), after we announce the index 25, User 1 with $\bm{\alpha}$ performs Picking Up, User 5 with $\bm{\lambda}$ performs Walking, and User 6 with $\bm{\beta}$ performs Sitting Down. Note that the scripts are put away before activities to minimize their impacts on data collection. Between the collection of each two samples, we first stop the collection of previous sample and then announce a new index to users, after which we start the collection of next sample. Therefore, we can ensure the purity of each sample, not containing nearby activities. More details of activities, users, user groups, and simultaneous scripts are provided in Appendix A.
Figure 3: Scripts that instruct users to perform identical/different activities at varying locations independently yet simultaneously.
Data Annotation. Since we apply the aforementioned scripts to instruct user activities, we have actually created data annotations before collection. We label each sample as “act_$<$group$>$_$<$sample$>$”, where “$<$group$>$” is the user group index, and “$<$sample$>$” is the sample index. For example, Fig. 0(a) shows the sample labeled as “act_30_25”, manifesting that it is the 25th sample of the 30th user group. We associate each label with the environment, WiFi band, number of users, identities, locations, and activities, to enable different sensing tasks. Note that the videos share the same labels with CSI and thereby can act as references to support unexplored tasks (e.g., multi-user pose estimation).